OpenStudy (anonymous):

Solve the following proportion for v 7/v = 3/13 Round your answer to the nearest tenth.

OpenStudy (anonymous):

@Godlovesme @Directrix @peterg2001

OpenStudy (alexandervonhumboldt2):

v=7*13/3

OpenStudy (alexandervonhumboldt2):

v=91/3=30 1/3